# Truth in the Few: High-Value Data Selection for Efficient Multi-Modal Reasoning

### v1

@@ 4.3 Key Insights and Discussion
Why does less data outperform the full dataset? The above empirical analysis shows that models trained on carefully curated data can notably outperform those trained on the entire dataset. To further elucidate the underlying mechanisms driving this “truth in the few” phenomenon, we examine sample difficulty distributions quantified by group-wise response accuracy in the GRPO generation paradigm in Figure 7(b). The results reveal that cognitive samples contain significantly fewer easy samples compared to the full dataset. Therefore, under conventional large-scale training, which usually restricts training to 1 or 2 epochs, the models lack sufficient repeated exposure to challenging samples, thus limiting the improvements brought from these valuable instances.
@@
Moreover, to further validate our hypothesis, we conduct a comparison in Figure 7(c), where models are trained on equivalent numbers of easy, hard, and cognitive samples. The results reveal that the abundance of easy examples contributes little to advancing the model’s reasoning capability and predominantly introduces redundant information. While challenging samples yield better performance than easy samples, they still fall short compared to the cognitive samples selected through our RAP method. We argue that the superiority of cognitive samples arises not merely from sample difficulty but from their ability to activate the model’s multi-modal reasoning capacity. In multi-modal contexts, the training samples must facilitate the activation of the model’s multi-modal reasoning mechanisms.
@@
Discussion on potential optimization during RL training. The static nature of our current RAP method motivates us to explore potential optimizations through the dynamic adjustment of the training dataset during RL training. Specifically, we examine the evolving distribution of cognitive samples identified by RAP criteria across successive epochs. Initially, as depicted in Figure 7(d), approximately 5,000 samples meet the cognitive sample criteria. However, the number of samples that continues to satisfy this criterion progressively declines as training advances. This observed trend suggests the promise of adaptive data sampling, wherein cognitive samples are re-sampled dynamically from the entire dataset after the first and third training epochs. Preliminary experiments indicate that such adaptive strategies yield tangible improvements in model performance. However, despite these advantages, the re-screening of the entire dataset after each epoch incurs a significant computational overhead. As a result, our final approach strategically refrains from adopting this adaptive data selection in order to maintain computational efficiency. This highlights the inherent trade-off between performance enhancements and computational cost when incorporating dynamic data selection strategies into the RL training process.
5 Conclusion

### v2

@@ IV-C Key Insights and Discussion
Why does less data outperform the full dataset? The above empirical analysis shows that models trained on carefully curated data can notably outperform those trained on the entire dataset. To further elucidate the underlying mechanisms driving this “truth in the few” phenomenon, we examine sample difficulty distributions quantified by group-wise response accuracy in the GRPO generation paradigm in Fig. 7(b). The results reveal that cognitive samples contain significantly fewer easy samples compared to the full dataset. Therefore, under conventional large-scale training, which usually restricts training to 1 or 2 epochs, the models lack sufficient repeated exposure to challenging samples, thus limiting the improvements brought from these valuable instances.
@@
Moreover, to further validate our hypothesis, we conduct a comparison in Fig. 7(c), where models are trained on equivalent numbers of easy, hard, and cognitive samples. The results reveal that the abundance of easy examples contributes little to advancing the model’s reasoning capability and predominantly introduces redundant information. While challenging samples yield better performance than easy samples, they still fall short compared to the cognitive samples selected through our RAP method. We argue that the superiority of cognitive samples arises not merely from sample difficulty but from their ability to activate the model’s multi-modal reasoning capacity. In multi-modal contexts, the training samples must facilitate the activation of the model’s multi-modal reasoning mechanisms.
TABLE VI: Performance comparison on two text-only benchmarks MATH500 [54] and AMC23 using Qwen2.5-7B [1] and LLaMA3.1-8b [55], with GRPO as the RL algorithm. Random and Full refer to training with randomly sampled data and the complete dataset, respectively.
Method MATH500 AMC23
Qwen2.5-7b-Random 72.40 50.25
Qwen2.5-7b-Full 74.40 51.25
Qwen2.5-7b-RAP (Ours) 74.60 51.75
LLaMA3.1-8b-Random 46.60 20.50
LLaMA3.1-8b-Full 48.60 22.75
LLaMA3.1-8b-RAP (Ours) 49.90 23.25
Generalization on text-only scenarios. Our RAP method is primarily designed to enhance multi-modal reasoning during RL post-training, and thus its core CDE relies on cross-modal signals, which is not directly transferable to unimodal settings. To address potential concerns regarding applicability to LLMs, we adapt the ACE and DRM modules to text-only scenarios and evaluate them on two text-only benchmarks, MATH500 [54] and AMC23, using the GRPO [3] as the RL algorithm. We conduct experiments on two representative LLMs, Qwen2.5-7B [1] and LLaMA3.1-8B [55]. In this adaptation, due to the absence of CDE, we implement DRM by directly removing easy samples defined by $D_{\text{iff}}=0$, i.e., samples consistently answered correctly across all trials, which typically provide limited optimization signal. As shown in Table VI, ACE and DRM still yield consistent gains over Random and Full training across both models, indicating that these components capture a generally beneficial data selection principle beyond multi-modal settings. However, the improvements are smaller than those in the multi-modal scenario. We attribute this mainly to two reasons: (1) RAP is originally motivated and optimized for multi-modal contexts where cross-modal discrepancies provide stronger supervision; (2) without CDE, DRM reduces to a difficulty-based filtering strategy, which is less targeted than the original design, thereby limiting the achievable gains.
@@
Discussion on potential optimization during RL training. The static nature of our current RAP method motivates us to explore potential optimizations through the dynamic adjustment of the training dataset during RL training. Specifically, we examine the evolving distribution of cognitive samples identified by RAP criteria across successive epochs. Initially, as depicted in Fig. 7(d), approximately 5,000 samples meet the cognitive sample criteria. However, the number of samples that continues to satisfy this criterion progressively declines as training advances. This observed trend suggests the promise of adaptive data sampling, wherein cognitive samples are re-sampled dynamically from the entire dataset after the first and third training epochs. Preliminary experiments indicate that such adaptive strategies yield tangible improvements in model performance. However, despite these advantages, the re-screening of the entire dataset after each epoch incurs a significant computational overhead. As a result, our final approach strategically refrains from adopting this adaptive data selection in order to maintain computational efficiency. This highlights the inherent trade-off between performance enhancements and computational cost when incorporating dynamic data selection strategies into the RL training process.
V Conclusion